# MU-Bench: A Multitask Multimodal Benchmark for Machine Unlearning

### v1

@@ Abstract
Recent advancements in Machine Unlearning (MU) have introduced solutions to selectively remove certain training samples, such as those with outdated or sensitive information, from trained models. Despite these advancements, evaluation of MU methods have been inconsistent, employing different trained models and architectures, and sample removal strategies, which hampers accurate comparison. In addition, prior MU approaches have mainly focused on singular tasks or modalities, which is not comprehensive. To address these limitations, we develop MU-Bench, the first comprehensive benchmark for MU that (i) unifies the sets of deleted samples and trained models, and (ii) provides broad coverage of tasks and data modalities, including previously unexplored domains such as speech and video classification. Our evaluation show that RandLabel (Graves et al., 2021) and SalUn (Fan et al., 2024b) are the most effective general MU approaches on MU-Bench, and Bad-T (Chundawat et al., 2023) and SCRUB (Kurmanji et al., 2023) are capable of achieving random performance on the deletion set. We analyze several under-investigated aspects of unlearning, including scalability, the impacts of parameter-efficient fine-tuning and curriculum learning, and susceptibility to dataset biases. MU-Bench provides an easy-to-use package that includes dataset splits, models, and implementations, together with a leader board to enable unified and scalable MU research [^1] Code: https://github.com/CLU-UML/MU-Bench.
@@ 2.1 Problem Formulation
Let $D_{\mathrm{Train}}$ denotes the training dataset, $D_{f}\subseteq D_{\mathrm{Train}}$ the subset to be unlearned, and $D_{r}=D_{\mathrm{Train}}\textbackslash D_{f}$ the remaining dataset post-unlearning. Given a model $f$ trained on $D_{\mathrm{Train}}$, machine unlearning seeks to remove the influence of $D_{f}$ from $f$ without affecting the knowledge it gained from $D_{r}$, without retraining from scratch. We term $f$ as the original model and $f^{\prime}$ as the model post-unlearning. A successful unlearned model $f^{\prime}$ should be minimally impacted by $D_{f}$, while maintaining the performance of $f$ on the original downstream test set $D_{\mathrm{Test}}$.
@@ 2.3 Unified Unlearning
We include the datasets, standardized data splits, evaluation scripts, and unlearning methods within an easy-to-use Python package and integerate them with commonly-used packages such as PyTorch (Paszke et al., 2019), Huggingface Transformers (Wolf et al., 2020), and Diffusers (von Platen et al., 2022), containing pretrained diffusion models for image and speech data. Users can initiate an unlearning experiment with minimal adjustment to existing script. All original model checkpoints are released for standardized unlearning and fair comparisons. We also host and maintain a leaderboard to rank methods overall and on individual tasks and architectures. For example, to remove 5% of training data from a BERT-base model trained on IMDB using Bad-T (Chundawat et al., 2023), only a minimal script modification is required. This setup simplifies the unlearning process and enables rapid comparison against methods and architectures.
1from transformers import TrainingArguments, AutoTokenizer, AutoModelForSequenceClassification
2
3# Additional code for unlearning
4from benchmark import UnlearningTrainer, UnlearningArguments
5unlearn_config = UnlearningArguments(
6 unlearn_method="bad_teaching", # MU method
7 backbone="bert-base", # Network architecture
8 data_name="imdb", # Dataset
9 del_ratio=5 # Standardized splits
10)
11
12# Standard HuggingFace code
13args = TrainingArguments(output_dir="tmp")
14original_model_name = f"mu-bench/{unlearn_config.backbone}-{unlearn_config.data_name}"
15raw_datasets = load_dataset(unlearn_config.data_name)
16tokenizer = AutoTokenizer.from_pretrained(original_model_name)
17original_model = AutoModelForSequenceClassification.from_pretrained(original_model_name)
18
19# Replace original HF trainer with our new trainer
20trainer = UnlearningTrainer(model=original_model, args=args, unlearn_config=unlearn_config, raw_datasets=raw_datasets, tokenizer=tokenizer)
21
22# Start unlearning and evaluation
23trainer.unlearn()
Listing 1: Example usage of MU-Bench: deleting 5% data from BERT-base trained on IMDB.
@@
Table 2: Taxonomy of unlearning techniques. Despite different formulations and loss functions, existing approaches can be viewed in a unified teacher-student framework, with three design choices: (i) knowledge measurement (KM), (ii) knowledge corruption on $D_{f}$ (Corrupt), and (iii) knowledge retention on $D_{r}$ (Retain). The combination of (i) and (ii) leads to teacher on $D_{f}$, while combination of (i) and (iii) leads to teacher on $D_{r}$. For teachers on $D_{f}$ and $D_{r}$, Loss represents the expected task loss $\mathbb{E}_{(x,y)\in D}\sum L(f(x),y)$ on $D_{f}$ and $D_{r}$. Rep. denotes the KL Divergence of output distribution $\mathbb{E}_{(x,y)\in D}\sum\mathrm{KL}(f^{\prime}(x),f(x))$ on $D_{f}$ and $D_{r}$. Trainable parameters are denoted as Dense or Sparse, and Internal or External. A detailed version of this table is available in Appendix LABEL:xx.
@@ 3 Experiments
Figure 2: Overall average accuracy across all discriminative tasks.
@@ 3.1 Main Results on Discriminative Tasks
As Figure 2 illustrates, NegGrad typically results in low performance on $D_{f}$, but severely compromises the knowledge on $D_{\mathrm{Test}}$ and $D_{r}$, indicating it is not an effective MU approach. In general, tasks like audio classification, video classification, text summarization and generation consistently challenge existing MU algorithms, potentially due to strong correlations within the data, see Figure 11-12). We report the average performance across all tasks as all metrics range from 0% to 100%.
@@ 4 Discussion and Analysis
Which design choices are effective for machine unlearning?
Is a lower performance on $D_{f}$ always better?
@@ A.1 Related work
Unlearning for generation models centers on removing copyrighted, private, NSFW, or biased content from generative models, including diffusion models (Gandikota et al., 2023; Kumari_2023_ICCV; Gandikota et al., 2024; Liu et al., 2024; Fuchi & Takagi, 2024; Fan et al., 2024b), image-to-image models (Li et al., 2024a), text summarization models (Chen & Yang, 2023), translation models (Wang et al., 2023), and text generation models (Lu et al., 2022; Jang et al., 2023; Kassem et al., 2023; Chen & Yang, 2023).

### v2

@@ Abstract
Recent advancements in Machine Unlearning (MU) have introduced solutions to selectively remove certain training samples, such as those with outdated or sensitive information, from trained models. Despite these advancements, evaluation of MU methods have been inconsistent, employing different trained models and architectures, and sample removal strategies, which hampers accurate comparison. In addition, prior MU approaches have mainly focused on singular tasks or modalities, which is not comprehensive. To address these limitations, we develop MU-Bench, the first comprehensive benchmark for MU that (i) unifies the sets of deleted samples and trained models, and (ii) provides broad coverage of tasks and data modalities, including previously unexplored domains such as speech and video classification. Our evaluation show that RandLabel (Graves et al., 2021) and SalUn (Fan et al., 2024b) are the most effective general MU approaches on MU-Bench, and Bad-T (Chundawat et al., 2023) and SCRUB (Kurmanji et al., 2023) are capable of achieving random performance on the deletion set. We analyze several under-investigated aspects of unlearning, including scalability, the impacts of parameter-efficient fine-tuning and curriculum learning, and susceptibility to dataset biases. MU-Bench provides an easy-to-use package that includes dataset splits, models, and implementations, together with a leader board to enable unified and scalable MU research. [^1] Project page: https://clu-uml.github.io/MU-Bench-Project-Page..
@@ 2.1 Problem Formulation
Let $D_{\mathrm{Train}}$ denotes the training dataset, $D_{f}\subseteq D_{\mathrm{Train}}$ the subset to be unlearned, and $D_{r}=D_{\mathrm{Train}}\backslash D_{f}$ the remaining dataset post-unlearning. Given a model $f$ trained on $D_{\mathrm{Train}}$, machine unlearning seeks to remove the influence of $D_{f}$ from $f$ without affecting the knowledge it gained from $D_{r}$, without retraining from scratch. We term $f$ as the original model and $f^{\prime}$ as the model post-unlearning. A successful unlearned model $f^{\prime}$ should be minimally impacted by $D_{f}$, while maintaining the performance of $f$ on the original downstream test set $D_{\mathrm{Test}}$.
@@ 2.2 Datasets and Tasks
1# Standard HuggingFace code
2from transformers import TrainingArguments
3args = TrainingArguments(output_dir="tmp")
4
5# Additional code for unlearning
6from mubench import UnlearningArguments, unlearn_trainer
7unlearn_config = UnlearningArguments(
8 unlearn_method="multi_delete", # MU method, MultiDelete ECCV’24
9 backbone="vilt", # Network architecture
10 data_name="nlvr2", # Dataset
11 del_ratio=5 # Standardized splits
12)
13trainer = unlearn_trainer(unlearn_config.unlearn_method)(
14 args=args,
15 unlearn_config=unlearn_config
16)
17trainer.unlearn() # Start Unlearning and Evaluation!
Listing 1: Example usage of MU-Bench: Unlearning 5% data from the VILT model trained on NLVR2 using MultiDelete.
@@ 2.3 Unified Unlearning
We include the datasets, standardized data splits, evaluation scripts, and unlearning methods within an easy-to-use Python package and integrate them with commonly-used packages such as PyTorch (Paszke et al., 2019), Huggingface Transformers (Wolf et al., 2020), and Diffusers (von Platen et al., 2022), containing pre-trained diffusion models for image and speech data. Users can initiate an unlearning experiment with minimal adjustment to existing script. All original model checkpoints are released for standardized unlearning and fair comparisons. We also host and maintain a leaderboard to rank methods overall and on individual tasks and architectures. For example, to remove 5% of training data from a BERT-base model trained on IMDB using Bad-T (Chundawat et al., 2023), only a minimal script modification is required shown in code example 1. This setup simplifies the unlearning process and enables rapid comparison against methods and architectures.
@@
Table 2: Taxonomy of unlearning techniques. Despite different formulations and loss functions, existing approaches can be viewed in a unified teacher-student framework, with three design choices: (i) knowledge measurement (KM), (ii) knowledge corruption on $D_{f}$ (Corrupt), and (iii) knowledge retention on $D_{r}$ (Retain). The combination of (i) and (ii) leads to teacher on $D_{f}$, while combination of (i) and (iii) leads to teacher on $D_{r}$. For teachers on $D_{f}$ and $D_{r}$, Loss represents the expected task loss $\mathbb{E}_{(x,y)\in D}\sum L(f(x),y)$ on $D_{f}$ and $D_{r}$. Rep. denotes the KL Divergence of output distribution $\mathbb{E}_{(x,y)\in D}\sum\mathrm{KL}(f^{\prime}(x),f(x))$ on $D_{f}$ and $D_{r}$. Trainable parameters are denoted as Dense or Sparse, and Internal or External.
@@ 3.1 Main Results on Discriminative Tasks
As Figure 2 illustrates, NegGrad typically results in low performance on $D_{f}$, but severely compromises the knowledge on $D_{\mathrm{Test}}$ and $D_{r}$, indicating it is not an effective MU approach. In general, tasks like audio classification, video classification, text summarization and generation consistently challenge existing MU algorithms, potentially due to strong correlations within the data, see Figure 11-12). We report the average performance across all tasks as all metrics range from 0 to 100%.
@@
Figure 2: Overall average accuracy across all discriminative tasks.
@@ 4 Discussion and Analysis
Which design choices are effective for machine unlearning?
Is a lower performance on $D_{f}$ always better?
@@ A.1 Related work
Unlearning for generation models centers on removing copyrighted, private, NSFW, or biased content from generative models, including diffusion models (Gandikota et al., 2023; Kumari et al., 2023; Gandikota et al., 2024; Liu et al., 2024; Fuchi & Takagi, 2024; Fan et al., 2024b), image-to-image models (Li et al., 2024a), text summarization models (Chen & Yang, 2023), translation models (Wang et al., 2023), and text generation models (Lu et al., 2022; Jang et al., 2023; Kassem et al., 2023; Chen & Yang, 2023).